Tutors Your help is needed please: Thank you very much
Solve
log2(x+5)+log2(x-1)=4
The 2s are smaller and below log

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(21583)
Answer by stanbon(48535)

Solve
log2(x+5)+log2(x-1)=4
---------------
log2[(x+5)(x-1)] = 4
[(x+5)(x-1)] = 2^4
x^2+4x-5 = 16
x^2 + 4x -21 = 0
(x+7)(x-3) = 0
Positive solution:
x = 3
-----------------
Cheers,
Stan H.